OpenStudy (anonymous):

please help i give out medals :) write an equation for a line perpendicular to this: y-6 = 2(x-4)

OpenStudy (anonymous):

y-6 = 2(x-4) y=2x-8+6 y=2x-2 let x=0=>y=-2 from the equation we could say the slope is 2 the slope of the required line is -1/2 so the equation is y=-1/2(X) +c substituting the points to find c,,,,, we get c = -2 so , the equation is y=-1/2(X)+ c

OpenStudy (anonymous):

y=2x-8+6 y=2x+2--- in slope form( y=mxb). to write a perpendicular eq. change the sign and and the reciprocal of the slope (mx) so, is y=- 1/2x+2 the answer

OpenStudy (anonymous):

oh and before what i did with the original equations y-6 = 2(x-4) is i think is in y-intercept form, i dont remember really its been years since i haven't seen this, you solve to get y=mx+b, m being the slope and b being the y-intercept in the graph. so, solve the parenthesis first, your left with y-6=2x-8. then add 6 to both sides, so it cancels out in the left and your left with y=2x-2. and then u do what i told you in the last post... hope it helps :)